# CRoP: Context-wise Robust Static Human-Sensing Personalization

### v4

@@ Abstract.
The advancement in deep learning and internet-of-things have led to diverse human sensing applications. However, distinct patterns in human sensing, influenced by various factors or contexts, challenge the generic neural network model’s performance due to natural distribution shifts. To address this, personalization tailors models to individual users. Yet most personalization studies overlook intra-user heterogeneity across contexts in sensory data, limiting intra-user generalizability. This limitation is especially critical in clinical applications, where limited data availability hampers both generalizability and personalization. Notably, intra-user sensing attributes are expected to change due to external factors such as treatment progression, further complicating the challenges. To address the intra-user generalization challenge, this work introduces CRoP, a novel static personalization approach. CRoP leverages off-the-shelf pre-trained models as generic starting points and captures user-specific traits through adaptive pruning on a minimal sub-network while preserving generic knowledge in the remaining parameters. CRoP demonstrates superior personalization effectiveness and intra-user robustness across four human-sensing datasets, including two from real-world health domains, underscoring its practical and social impact. Additionally, to support CRoP’s generalization ability and design choices, we provide empirical justification through gradient inner product analysis, ablation studies, and comparisons against state-of-the-art baselines.
Intra-user Generalization, Personalization, Contest-wise Robustness
@@ 1. Introduction
Recent automated human sensing applications—like activity recognition, fall detection, and health tracking - revolutionize daily life, especially in personal health management (Wang et al., 2023a). However, unique user patterns and natural distribution shifts (Gong et al., 2023) caused by behaviors, physical traits, environment, and device placements (Ustev et al., 2013; Stisen et al., 2015) lead to the underperformance of generic sensing models in practical use. To tackle this, various domain adaptation techniques have been explored, with personalization widely used to adapt a generic model to the target user’s specific domain or natural distribution (Lamichhane et al., 2023; Iaboni et al., 2022; Meegahapola et al., 2023; Ahamed and Farid, 2018; Ren et al., 2022; Sempionatto et al., 2021; Boukhechba et al., 2020). In literature, personalization occurs either during the enrollment phase (static) (Duan et al., 2023; Liu et al., 2022; Burns et al., 2022) or continuously throughout application use, a process known as continual learning (Daniels et al., 2023; Liu et al., 2024; Wu et al., 2024; Wang et al., 2022a).
Continual learning methods involve retraining models with new data, either supervised or unsupervised. While these approaches enable models to adapt to new patterns and changes in data distribution over time, they often face efficiency challenges. As Harun et al. (2023) points out, these methods, particularly those that prevent catastrophic forgetting (Kirkpatrick et al., 2017; Wang et al., 2022a; Daniels et al., 2023; Fini et al., 2022; Tang et al., 2024a, b; Mallya and Lazebnik, 2018; Mallya et al., 2018; Wang et al., 2022b), often struggle with memory, computation, and storage requirements inefficiencies, limiting their real-world applicability. Frequent model updates on user devices can introduce significant delays. This is especially true for devices with limited processing power, such as smartphones or wearables. These delays reduce the responsiveness of the application. They also lead to battery drain and inefficiency in real-time applications. Furthermore, to avoid catastrophic forgetting, several continual learning approaches rely on replay-based methods which require storage of previously encountered data (van de Ven and Tolias, 2019; Hayes et al., 2019; van de Ven et al., 2020). This not only increases storage requirements but also raises privacy concerns. Constantly storing potentially identifiable information increases the risk of data breaches or misuse. This is especially concerning in regulatory environments, such as healthcare (Verma et al., 2023).
@@
In contrast, static personalization offers an efficient alternative by customizing the model with a one-time, limited dataset collected during enrollment. This approach minimizes computation, requires no ongoing data or label storage or collection, and reduces user engagement, making it especially suitable for resource-constrained human-sensing applications. However, existing such studies often overlook intra-user variability due to factors like changes in magnetic field (Robert-Lachaine et al., 2017), sensor position (Park et al., 2014), terrain (Kowalsky et al., 2021), or the health symptoms (Päeske et al., 2023), leading to poor intra-user generalizability for contexts not present during personalization. For instance, a smartphone activity recognition model personalized with handheld data may perform poorly when the phone is in a pocket.
Static personalization is particularly crucial for clinical datasets, which often suffer from data scarcity, leading to reduced robustness of lab-validated models for prospectively collected users (Berisha et al., 2021). It enhances model accuracy for clinical users whose traits are underrepresented in the global model’s training data. In contrast, continuous supervised personalization is generally infeasible in many health domains since ground truths must be validated by clinicians, making it impractical in continuous settings, especially in remote or mobile health applications.
Notably, the distribution of clinical data is expected to shift, even within the same individual. For instance, in clinical speech technologies, changes in data distribution over time may occur due to the progression of neurodegenerative diseases, relevant for disease monitoring apps (Stegmann et al., 2020), or through desired learning mechanisms resulting from the use of technology, as seen in automated speech therapy apps (Benway and Preston, 2023). Similarly, in wearable-based stress monitoring, psychophysiological data distribution varies as individuals face different stressors (Nagaraj et al., 2023). This research defines ‘context’ as the intra-user data distribution formed by varying factors.
This research gap is worsened since static personalization typically relies on a small sample set from the target user, covering limited contexts—particularly in clinical settings or applications with data scarcity (Berisha et al., 2021; Benway and Preston, 2023). Commercial human sensing technologies like Google Assistant, Amazon Alexa, and Apple’s Siri also statically personalize speech recognition models using limited phrases (Team, 2017; Awobajo, 2023; Phelan, 2019). Similarly, the Apple Watch uses initial calibration for enhanced running activity tracking (Apple, 2023; Potuck, 2021). This limited context during personalization is problematic, as shown in Section 3, where we demonstrate that static personalization may improve performance in training contexts but can also significantly degrade it in other unseen contexts for the same user.
Figure 1. Problem Setting
@@
In summary, Figure 1 outlines the paper’s scope and objective. The goal is to develop an intra-user robust approach to personalize an off-the-shelf generic model for a specific user using limited data from limited contexts. The primary objective is to ensure that the personalized model thus obtained exhibits robust generalization capabilities across unseen contexts. Crucially, unseen context data takes no part in training or adjusting the personalized model outcome, and the generic model remains entirely off-the-shelf, with no accessibility for modification or design choices. These constraints highlight the real-world impact of this research, particularly in clinical settings where privacy concerns often limit data sharing (Malin et al., 2018; Rathbone et al., 2023), and only trained off-the-shelf models are shared among researchers and developers.
@@
    Personalization with intra-user generalization has two challenges: i) learning user-specific patterns and ii) keeping information about generic patterns intact. During personalization, CRoP leverages pruning and regularization to capture the user-specific patterns present in the available-context data on a minimal sub-network of the model. The remaining parameters of the model are utilized to retain generic patterns that are not present in the available context, i.e., personalization data, thus achieving intra-user generalizability to unseen contexts. Pruning is widely used in other areas, including in continuous learning literature (e.g., Packnet (Mallya and Lazebnik, 2018), and Piggyback (Mallya et al., 2018)), to obtain sub-networks for each new task which not only requires data from every task (here unseen context) but also needs identification of the task differences (i.e., difference among unseen contexts) in order to select the sub-network for inference. In contrast, this is the first paper to use pruning with adaptive intensity to capture the user’s traits in a compressed sub-network, thus striking a balance between personalization and robustness in new, unseen contexts for the same user without identifying the contexts, their differences, or requiring data from unseen contexts for adaptation.
@@
    To showcase CRoP’s efficacy, comprehensive evaluations were performed on four human sensing datasets: PERCERT-R (Benway et al., 2023): a clinical speech therapy dataset, WIDAR (Zhang et al., 2022): a lab-based WiFi-CSI dataset, ExtraSensory: a real-world mobile sensing dataset (Vaizman et al., 2017), and a stress-sensing dataset (Xiao et al., 2024), while considering two disjoint contexts for each dataset. In order to obtain information about context variation within PERCEP-R and Stress-sensing, we collaborated with the original authors of these two datasets. This additional annotation will be released alongside this publication.
@@
    On an average across all datasets, CRoP shows a personalization benefit of 35.23 percent points and generalization benefit of 7.78 percent points. As compared to the best baseline (Packnet), these gains are 9.18 and 9.17 percent points higher, respectively. Moreover, alongside a detailed ablation study discussion in Section 8, an empirical justification of CRoP’s design choices that enable intra-user generalizability among different contexts is provided in Section 7.1, employing Gradient Inner Product(GIP) (Shi et al., 2021) analysis. Additionally, in order to demonstrate the feasibility of on-device personalization through CRoP, we compute training time and resource requirements for five platforms or devices.
The paper is arranged in different sections. Section 2 reviews the related work, while Section 3 presents the results of a preliminary study on the WIDAR datasets, which motivates the proposed approach. Section 4 provides a formal description of the problem statement, followed by Section 5, which outlines the detailed methodology. Section 6 covers the experimental setup, and results. In Section 7, we present an empirical study to justify the approach, and Section 8 offers an ablation study to support various design choices. Section 9 discusses runtime analysis, Section 10 explores limitations and future directions, and Section 11 addresses the broader impact of the work. Finally, Section 12 concludes the study.
The work is accompanied by an extensive appendix, which provides details of experiment setup, hyperparameters, and link to code ensuring reproducibility in Appendix A. Additionally, Appendix B provides a detailed analysis of the user-specific results for each dataset along with a detailed ablation study.
@@ 2. Related Work
As discussed above, this work aims to personalize off-the-shelf models while ensuring intra-user generalizability leveraging limited-context data. While existing approaches attempt to address these adaptation challenges, they often impose restrictive conditions. Many require specialized training of the generic model with access to the generic data, assume knowledge of the target domain, or necessitate target users to share their data from a new/previous-unseen context for adapting the model, in some studies with labels or annotations. Additionally, some methods rely on repetitive model retraining, which hinders their suitability for real-world applications due to privacy concerns, computational inefficiency, or excessive user engagement.
@@ 2.1. Domain Adaptation and Generalization
Domain adaptation (DA) and Domain Generalization (DG) techniques address performance drops due to distributional shifts between source and target domains (Kouw and Loog, 2019). These methods are useful when models trained on generic data need to adapt to new users or contexts. The key difference is that while DA approaches have access to target domain data, DG techniques work without any target domain data, not even unlabeled (Wang et al., 2023b).
The domain generalization techniques can be categorized into Data Manipulation, Representation Learning, and other learning strategies (Wang et al., 2023b). Data Manipulation approaches include techniques like data augmentation and synthetic data generation, which are used to increase data diversity, helping models learn invariant features across different domains (Shankar et al., 2018; Volpi et al., 2018; Zhou et al., 2020). Representation Learning aims to capture domain-invariant features using techniques like adversarial training and embedding learning to improve model robustness, such as Invariant risk minimization (IRM)(Arjovsky et al., 2020) and domain-adversarial neural networks (DANN)(Ganin et al., 2016). Learning Strategies such as meta-learning (Li et al., 2018) and ensemble learning (D’Innocente and Caputo, 2019) train the model on multiple simulated domains to improve its ability to generalize. However, all of these approaches aim to train the generic model while this paper’s scope does not allow access to generic data and uses a pre-trained off-the-shelf model for adaptation.
Similarly, some domain adaptation approaches rely on generic model training while utilizing data from the target domain (Long et al., 2016). On the other hand, certain domain adaptation approaches require access to the source data used to train the generic model without requiring specialized training of the generic model. These approaches rely on discrepancy minimization and self-supervision to align source and target distributions by learning domain-invariant representations (Li et al., 2021; Ganin and Lempitsky, 2015). A few approaches that do not require access to the source domain data or generic model training are referred to as source-free domain adaptation (SFDA)(Liang et al., 2020). Liang et al. (2020) (SHOT) proposed the transfer of hypothesis from source by freezing the parameter weights for the classifier layers and only allowing the feature extraction layer to be finetuned to the new domain. The approach is applicable to unsupervised domain adaptation scenarios and employs self-supervised pseudo-labeling to align the target domain’s representations to the source hypothesis.
Given that the problem addressed by SHOT aligns with the criteria of having no access to generic data and avoiding specialized generic model training, we adopted SHOT as one of the baseline approaches in this study. However, it is important to note that SHOT does not effectively tackle the limitations posed by restricted-context data during the fine-tuning process. As a result, its performance may not adequately generalize to the intra-user variability present in the data, which is a crucial aspect of our research focus.
@@ 2.2. Continual Learning
Continual personalization approaches (Wang et al., 2022a; Daniels et al., 2023; Fini et al., 2022; Tang et al., 2024a, b; Mallya and Lazebnik, 2018; Mallya et al., 2018; Wang et al., 2022b) can improve intra-user generalizability by continually fine-tuning the model as new data arrives. Some of these approaches (Fini et al., 2022; Tang et al., 2024a, b) require specialized training of the generic model, limiting the use of off-the-shelf pre-trained models. Others like PackNet (Mallya and Lazebnik, 2018) and Piggyback (Mallya et al., 2018) propose supervised methods that require continued steam of labeled data, limiting their application in health-care scenarios. However, all continuous learning approaches require repeated computation overhead to adjust the model outcome to new data (Prabhu et al., 2023), which can be infeasible in real-time applications, more so for scalable platforms like wearables (Schmidt et al., 2018), which is prominent for health sensing such as stress or fall detection. Daniels et al. (2023) proposes a continual training framework that is compatible with edge devices and requires less training effort, but it uses a different and smaller model architecture at the personalized devices for generating feature embeddings. Thus hindering the direct usage of off-the-shelf models and increasing the training effort. Nevertheless, since the problems addressed by Packnet, and Piggyback are the closest to the problem addressed in this study, we considered these approaches as baselines.
Packnet (Mallya and Lazebnik, 2018) and Piggyback (Mallya et al., 2018) adapt the generic model for a stream of continuously changing tasks. Packnet relies on finetuning, pruning and re-training the model for each new task. On the other hand, Piggyback employs pruning to learn a new binary mask for each new task which is then applied to a generic model in order to get task-specific results. However, this requires correct identification of the task before choosing the mask. For every new task, the same procedure is repeated, hoping that different sets of parameters will be important for different tasks. This aligns well with our initial study, where we identify that different contexts’ data focus on different sets of parameters.
Notably, all continual learning approaches, including PackNet and Piggyback, face a significant limitation: they necessitate access to data from the target domain, whether in a supervised or unsupervised format, in order to effectively adapt the model. While these approaches successfully address the challenge of catastrophic forgetting, they primarily focus on contexts that the model has previously encountered, leaving them ill-equipped to adapt to entirely unseen domains. In contrast, our research framework specifically excludes access to data from these unseen contexts during model adaptation, highlighting a critical gap in the capabilities of existing continual learning methodologies. Consequently, PackNet, Piggyback, and similar approaches lack the mechanisms necessary to adapt to novel domains, underscoring the need for innovative strategies that can operate without prior exposure to the target data.

### v5

@@ Abstract.
The advancement in deep learning and internet-of-things have led to diverse human sensing applications. However, distinct patterns in human sensing, influenced by various factors or contexts, challenge the generic neural network model’s performance due to natural distribution shifts. To address this, personalization tailors models to individual users. Yet most personalization studies overlook intra-user heterogeneity across contexts in sensory data, limiting intra-user generalizability. This limitation is especially critical in clinical applications, where limited data availability hampers both generalizability and personalization. Notably, intra-user sensing attributes are expected to change due to external factors such as treatment progression, further complicating the challenges. To address the intra-user generalization challenge, this work introduces CRoP, a novel static personalization approach. CRoP leverages off-the-shelf pre-trained models as generic starting points and captures user-specific traits through adaptive pruning on a minimal sub-network while allowing generic knowledge to be incorporated in remaining parameters. CRoP demonstrates superior personalization effectiveness and intra-user robustness across four human-sensing datasets, including two from real-world health domains, underscoring its practical and social impact. Additionally, to support CRoP’s generalization ability and design choices, we provide empirical justification through gradient inner product analysis, ablation studies, and comparisons against state-of-the-art baselines.
Intra-user Generalization, Personalization, Context-wise Robustness, Robust Personalization, Adaptive Personalization
ccs: Human-centered computing Ubiquitous and mobile computing theory, concepts and paradigms
@@ 1. Introduction
Recent automated human sensing applications—like activity recognition, fall detection, and health tracking - revolutionize daily life, especially in personal health management (Wang et al., 2023a). However, unique user patterns and natural distribution shifts (Gong et al., 2023) caused by behaviors, physical traits, environment, and device placements (Ustev et al., 2013; Stisen et al., 2015) lead to the underperformance of generic sensing models in practical use. To tackle this, various domain adaptation techniques have been explored, with personalization widely used to adapt a generic model to the target user’s specific domain (Meegahapola et al., 2023; Sempionatto et al., 2021). In literature, personalization occurs either during the enrollment phase (static) (Duan et al., 2023; Liu et al., 2022; Burns et al., 2022) or continuously throughout application use, a process known as continual learning (Daniels et al., 2023; Wu et al., 2024).
Continual learning methods involve retraining models with new data, either supervised or unsupervised. While these approaches enable models to adapt to new patterns and changes in data distribution over time, they often face efficiency challenges. As Harun et al. (2023) points out, these methods, particularly those that prevent catastrophic forgetting (Daniels et al., 2023; Fini et al., 2022; Tang et al., 2024a, b; Mallya and Lazebnik, 2018; Mallya et al., 2018; Wang et al., 2022), often struggle with memory, computation, and storage requirements inefficiencies, limiting their real-world applicability. Frequent model updates on user devices can introduce significant delays. This is especially true for devices with limited processing power, such as smartphones or wearables. These delays reduce the responsiveness of the application. They also lead to battery drain and inefficiency in real-time applications. Furthermore, to avoid catastrophic forgetting, several continual learning approaches rely on replay-based methods which require storage of previously encountered data (van de Ven and Tolias, 2019; Hayes et al., 2019; van de Ven et al., 2020). This not only increases storage requirements but also raises privacy concerns. Constantly storing potentially identifiable information increases the risk of data breaches or misuse. This is especially concerning in regulatory environments, such as healthcare (Verma et al., 2023).
@@
In contrast, static personalization offers an efficient alternative by customizing the model with a one-time, limited dataset collected during enrollment. This approach minimizes computation, requires no ongoing data or label storage or collection, and reduces user engagement, making it especially suitable for resource-constrained human-sensing applications. However, existing such studies often overlook intra-user variability due to factors like changes in magnetic field (Robert-Lachaine et al., 2017), sensor position (Park et al., 2014), terrain (Kowalsky et al., 2021), or the health symptoms (Päeske et al., 2023), leading to poor intra-user generalizability for contexts not present during personalization. For instance, a smartphone activity recognition model personalized with handheld data may perform poorly when the phone is in a pocket. Notably, the distribution of clinical data is expected to shift, even within the same individual. For instance, in clinical speech technologies, changes in data distribution over time may occur due to the progression of neurodegenerative diseases, relevant for disease monitoring apps (Stegmann et al., 2020), or through desired learning mechanisms resulting from the use of technology, as seen in automated speech therapy apps (Benway and Preston, 2023).
This research gap is worsened since static personalization typically relies on a small sample set from the target user, covering limited available contexts—particularly in clinical settings or applications with data scarcity (Berisha et al., 2021; Benway and Preston, 2023). Commercial human sensing technologies like Google Assistant, Amazon Alexa, and Apple’s Siri also statically personalize speech recognition models using limited phrases (Team, 2017; Awobajo, 2023; Phelan, 2019). Similarly, the Apple Watch uses initial calibration for enhanced running activity tracking (Apple, 2023; Potuck, 2021). This limited available context during personalization is problematic, as shown in Section 3, where we demonstrate that static personalization may improve performance in the available training contexts but can also significantly degrade it in other unseen contexts for the same user.
@@
Figure 1. Problem Setting: Study Objective
In summary, Figure 1 outlines the paper’s scope and objective. The goal is to develop an intra-user robust approach to personalize an off-the-shelf generic model for a specific user using limited data from limited available contexts. The primary objective is to ensure that the personalized model thus obtained exhibits robust generalization capabilities across all contexts (i.e., available and unseen contexts). Crucially, unseen context data takes no part in training or adjusting the personalized model outcome, and the generic model remains entirely off-the-shelf, with no accessibility for modification or design choices. These constraints highlight the real-world impact of this research, particularly in clinical settings where privacy concerns often limit data sharing (Malin et al., 2018; Rathbone et al., 2023), and only trained off-the-shelf models are shared among researchers and developers.
@@
    Personalization with intra-user generalization has two challenges: i) learning user-specific patterns and ii) keeping information about generic patterns intact. CRoP addresses these challenges by leveraging adaptive pruning and regularization followed by a mixing step to optimize this balance. After initial finetuning or model adaptation for the target user, CRoP (its ToleratedPrune module) dynamically adapts pruning levels per user to extract a personalized substructure that best represents the user’s traits in the available context. When users require significant parameter adjustments, CRoP’s regularizer increases parameter polarization, enabling more aggressive pruning and restoration of more parameter values, i.e., information from the generic model in the subsequent mixing step. This adaptive pruning strategy dynamically balances personalization and generalization by tailoring the recovery of generic information to user needs.
@@
    To showcase CRoP’s efficacy, comprehensive evaluations were performed on four human sensing datasets: PERCERT-R (Benway et al., 2023): a clinical speech therapy dataset, WIDAR (Zhang et al., 2022): a lab-based WiFi-CSI dataset, ExtraSensory: a real-world mobile sensing dataset (Vaizman et al., 2017), and a stress-sensing dataset (Xiao et al., 2024), while considering two disjoint contexts for each dataset. In order to obtain information about context variation within PERCEPT-R and Stress-sensing, we collaborated with the original authors of these two datasets. This additional annotation will be released alongside this publication.
@@
    On average across all datasets and corresponding all contexts, CRoP achieves a 35.23 percentage point improvement in personalization compared to the generic model, and a 7.78 percentage point improvement in generalization compared to the conventional-finetuned model (Hong et al., 2016). As compared to the best baseline (Packnet), these gains are 9.18 and 9.17 percent points higher, respectively. Moreover, alongside a detailed ablation study discussion in Section 8, an empirical justification of CRoP’s design choices that enable intra-user generalizability among different contexts is provided in Section 7.1, employing Gradient Inner Product(GIP) (Shi et al., 2021) analysis. Additionally, in order to demonstrate the feasibility of on-device personalization through CRoP, we compute training time and resource requirements for seven platforms or devices.
@@ 2.1. Domain Adaptation and Generalization
Domain adaptation (DA) and Domain Generalization (DG) techniques address performance drops due to distributional shifts between source and target domains (Kouw and Loog, 2019). These methods are useful when models trained on generic data need to adapt to new users or contexts. The key difference is that while DA approaches have access to target domain data, DG techniques work without any target domain data, not even unlabeled (Wang et al., 2023b). Thus, the scenario addressed by DA approaches is more similar to our problem setting.
Domain adaptation approaches typically rely on target domain data (Long et al., 2016) or require access to source data without specialized training of the generic model (Li et al., 2021; Ganin and Lempitsky, 2015). Source-free domain adaptation (SFDA) methods, such as SHOT (Liang et al., 2020), eliminate the need for source data or retraining the generic model. SHOT transfers the source hypothesis by freezing classifier weights and fine-tuning the feature extractor using self-supervised pseudo-labeling to align target representations.
Given that the problem addressed by SHOT aligns with the criteria of having no access to generic data and avoiding specialized generic model training, we adopted SHOT as one of the baseline approaches in this study. However, it is important to note that SHOT does not effectively tackle the limitations posed by restricted-context data during the fine-tuning process. SHOT employs pseudo-labels generated using a combination of k-means clustering and nearest neighbor classification on target domain’s data. In our setting, that target domain would be user’s available context data. Since, SHOT’s training approach does not consider futher changes in the target domain, SHOT would require another round of training using unseen domain’s data which is unavailable in current setting. As a result, its performance may not adequately generalize to the intra-user variability present in the data, which is a crucial aspect of our research focus.
@@ 2.2. Continual Learning
Continual personalization approaches (Daniels et al., 2023; Fini et al., 2022; Tang et al., 2024a, b; Mallya and Lazebnik, 2018; Mallya et al., 2018; Wang et al., 2022) can improve intra-user generalizability by continually fine-tuning the model as new data arrives. Some of these approaches (Fini et al., 2022; Tang et al., 2024a, b) require specialized training of the generic model, limiting the use of off-the-shelf pre-trained models, while others (Mallya and Lazebnik, 2018; Mallya et al., 2018) require continued stream of labeled data, limiting their application in health-care scenarios. However, all continuous learning approaches require repeated computation overhead to adjust the model outcome to new data (Prabhu et al., 2023), which can be infeasible in real-time applications, more so for resource-constrained platforms like wearables (Schmidt et al., 2018), which is prominent for health sensing such as stress or fall detection.
Packnet (Mallya and Lazebnik, 2018) and Piggyback (Mallya et al., 2018) adapt the generic model for a stream of continuously changing tasks. Packnet relies on finetuning, pruning and re-training the model for each new task. On the other hand, Piggyback employs pruning to learn a new binary mask for each new task which is then applied to a generic model in order to get task-specific results. However, this requires correct identification of the task before choosing the mask.
Notably, all continual learning approaches, including PackNet and Piggyback, require target domain data to adapt, limiting their ability to handle entirely unseen domains. While they mitigate catastrophic forgetting, they focus on previously encountered contexts. In contrast, our framework excludes unseen context data during adaptation, exposing a key limitation of existing methods.